OpenStudy (anonymous):

Find the exact value: (pi)/12 I believe the formula needed is a sum and difference formula but I'm not sure. (Trig problem)

OpenStudy (unklerhaukus):

in terms of sine and cosine/ ?

OpenStudy (anonymous):

pi=180(in trigonometry) pi/12=180/12=15 degree

OpenStudy (anonymous):

half angle formula, as it is half of \(\frac{\pi}{6}\)